The following information is available for Riverbed Corporation for 2019 (its first year of operations).

1. Excess of tax depreciation over book depreciation, $36,800. This $36,800 difference will reverse equally over the years 2020–2023.
2. Deferral, for book purposes, of $18,300 of rent received in advance. The rent will be recognized in 2020.
3. Pretax financial income, $320,500.
4. Tax rate for all years, 20%.

(a) Compute taxable income for 2019.

(b) Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2019. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

(c) Prepare the journal entry to record income tax expense, deferred income taxes, and income taxes payable for 2020, assuming taxable income of $299,000. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1
a
Pretax financial income 320500
Less: Excess of tax depreciation -36800
Add: Rent received in advance 18300
Taxable income for 2019 302000
b
Account titles and Explanation Debit Credit
Income tax expense 64100
Deferred Tax Asset 3660 =18300*20%
       Deferred Tax Liability 7360 =36800*20%
       Income tax payable 60400 =302000*20%
c
Account titles and Explanation Debit Credit
Income tax expense 61620
Deferred Tax Liability 1840 =7360/4
       Deferred Tax Asset 3660
       Income tax payable 59800 =299000*20%